OpenStudy (anonymous):

Find the standard form of the equation of the ellipse with the given characteristics vertices: (-9, -11) (-9,7) minor axis of length: 16

OpenStudy (anonymous):

I got (x+9)^2/81+(y+2)^2/64=1 Is that right?

OpenStudy (jdoe0001):

well, look at the vertices locations => (-9, -11) (-9,7) notice the "x" coordinate doesn't change the "y" coordinate does that means the ellipse major axis is vertical so, what's the distance between -11 and +7? well, 18, half that is "a" a = 9 the "b" component is given, minor axis is 16, half that is "b" b = 8 now the center of the ellipse is half-way between the 2 vertices so what's the midpoint of (-9, -11) (-9,7) ? well x = -9 from -11 going up "a" units, or 9 units, then -11+9 = 2 so the center is at (-9, 2)

OpenStudy (anonymous):

am i right though? I knew all that its always the center that gets me

OpenStudy (jdoe0001):

the center is half-way through the vertices, you can always just use the "midpoint equation" for that, so you have in this case 2 points

OpenStudy (jdoe0001):

I just noticed using the midpoint formula, center should be (-9, -2) dohhh -11 up 9 units, lands still on the 3rd quadrant heheh

OpenStudy (jdoe0001):

you can always just use the "midpoint equation" for that, so you have in this case 2 points $$ (-9, -11) (-9,7)\\ \left(\cfrac{x_2 + x_1}{2},\cfrac{y_2 + y_1}{2} \right)\\ \left(\cfrac{-9 + (-9)}{2},\cfrac{7 + (-11)}{2} \right)\\ \left(-9,-2 \right) $$